OpenStudy (notamathgenius):

Write an equation of the line that passes through the given point and is parallel to the graph of the given equation.

OpenStudy (notamathgenius):

(0, -5); y = 9x

OpenStudy (anonymous):

A parallel line will have the same slope, which is "9". Now, we use the point-slope formula for the equation of a line: y - y1 = m(x - x1) y - (-5) = 9(x - 0) y + 5 = 9x y = 9x - 5 slope-intercept form

OpenStudy (anonymous):

All good now, @Notamathgenius ?

OpenStudy (notamathgenius):

yes :)

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Notamathgenius